OpenStudy (anonymous):

Select all of the following that are ordered pairs of the given function. f(x) = 3 - 2x (-2, -1) (-1, 5) (0, 3) (1, 0) (2, -1)

OpenStudy (bahrom7893):

plug in each point to test ex: -1=3-2(-2) -1=3-4 -1=-1 so yea the first one is

OpenStudy (bahrom7893):

repeat for other points

OpenStudy (anonymous):

^ The first one isn't.

OpenStudy (bahrom7893):

oh u're right lol

OpenStudy (bahrom7893):

-2*-2=+4

OpenStudy (anonymous):

The 2nd, 3rd and last ones.